# Counting observables in stochastic excursions

### v1

@@ Abstract
Although there are many interesting counting observables to be explored in this model, we restrict ourselves to the cold current, entropy production, and dynamical activity. The cold current $\hat{J}_{c}$ is obtained by setting weights $\nu_{\ell xy}=\pm 1$ in Eq. (3) if the transition $x\xrightarrow{\ell}y$ removes (resp. adds) an excitation from (resp. into) the cold reservoir. Excursions can fully characterize the probability distribution of having bounces (i.e. a trajectory where the cold current is zero), in contrast with the steady state analysis supp . In Fig. 2(a) we show the covariance between the cold current $\hat{J}_{c}$ and the excursion duration $\hat{\mathcal{T}}$ as a function of the cold bath occupation $n_{c}$. Interestingly, longer excursions tend to have larger $\hat{J}_{c}$ in the cooling window, which is not necessarily the case when $\langle\hat{J}_{c}\rangle>0$, and a vanishing covariance pinpoints the value of $n_{c}$ where the cooling window begins, see End Matter. In Fig. 2(b) we show the distribution of the dynamical activity $\hat{\mathcal{A}}$, revealing that bounces dominate the dynamics. Notice that dynamical activity can only have positive values, with a lower bound of $2$ (the shortest excursion length is $2$ which is a bounce), and they are always even due to the specific model topology and allowed transitions in $B$. We go a step further and show in Fig. 2(c) the waiting-time distributions of the first transition within $B$ provided it does not return to $A$ (i.e. excluding bounces). Interestingly, the transition associated with the three-body coupling ($10\to 10$ 101) is dominant for shorter times, but is strongly suppressed afterward.
@@
In Figs. 2(d) and (e), we characterize the diffusion coefficient (11) in terms of its components, both for the entropy production $D_{\Sigma}$ and the dynamical activity $D_{\mathcal{A}}$. In the former, the fluctuations are strongly dominated by the variance term of $\hat{Q}=\hat{\Sigma}$; whereas in the latter there is a clear competition between different contributions, with $E(\hat{\mathcal{A}})^{2}$ being the most relevant term. This result sheds light on the very nature of the noise in FCS. $\hat{\Sigma}$ is an observable that can have positive and negative values related by fluctuation theorems, whereas $\hat{\mathcal{A}}$ is strictly positive. The stark contrast of behavior on the two observables shows the versatility of the excursion framework to characterize fluctuations.
@@
There are many interesting questions to be addressed as a follow-up to the work presented here. Initially, we point out that the excursions themselves may be correlated when region $A$ has more than one state, and such correlations may be a useful resource. The dynamics we considered here is classical, but the concept of excursions itself is general and could be employed in a quantum setting. Finally, exploring what excursions have to say about generic fluctuation theorems and uncertainty relations, as well as inference of counting observables, is subject to further work.
@@ Acknowledgements.
GF acknowledges useful discussions with Felipe Pereira-Alves and Jiheng Duan. PH was supported by the project INTER/FNRS/20/15074473 funded by F.R.S.-FNRS (Belgium) and FNR (Luxembourg). This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Award Number DE-SC0025516.

### v2

@@ Abstract
Although there are many interesting counting observables to be explored in this model, we restrict ourselves to the cold current, entropy production, and dynamical activity. The cold current $\hat{J}_{c}$ is obtained by setting weights $\nu_{\ell xy}=\pm 1$ in Eq. (3) if the transition $x\to_{\ell}y$ removes (resp. adds) an excitation from (resp. into) the cold reservoir. Excursions can fully characterize the probability distribution of trajectories having zero cold current, in contrast with the steady state analysis supp . In Fig. 2(a) we show the covariance between the cold current $\hat{J}_{c}$ and the excursion duration $\hat{\mathcal{T}}$ as a function of the cold bath occupation $n_{c}$. Interestingly, longer excursions tend to have larger $\hat{J}_{c}$ in the cooling window, which is not necessarily the case when $\langle\hat{J}_{c}\rangle>0$, and a vanishing covariance pinpoints the value of $n_{c}$ where the cooling window begins, see End Matter. In Fig. 2(b) we show the distribution of the dynamical activity $\hat{\mathcal{A}}$, revealing that shorter excursions dominate the dynamics. Note that dynamical activity is lower bounded by $2$ (which is the shortest excursion length), and they are always even due to the specific model topology and allowed transitions in $B$. We go a step further and show in Fig. 2(c) the waiting-time distributions of the first transition within $B$ provided it does not return to $A$. Interestingly, the transition associated with the three-body coupling ($10\to 10$ 101) is dominant for shorter times, but is strongly suppressed afterward.
@@
In Figs. 2(d) and (e), we characterize the diffusion coefficient (11) in terms of its components for the entropy production $D_{\Sigma}$ and the dynamical activity $D_{\mathcal{A}}$. In the former, the fluctuations are strongly dominated by the variance term of $\hat{\Sigma}$; whereas in the latter there is a clear competition between different contributions, with $E(\hat{\mathcal{A}})^{2}$ being the most relevant term. This result sheds light on the very nature of the noise in FCS. $\hat{\Sigma}$ is an observable that can have positive and negative values related by fluctuation theorems, whereas $\hat{\mathcal{A}}$ is strictly positive. The stark contrast of behavior on the two observables shows the versatility of the excursion framework to characterize fluctuations.
@@
There are many questions to be addressed as a follow-up to the work presented here. Initially, we point out that the excursions themselves may be correlated when region $A$ has more than one state, and such correlations may be a useful resource. We considered classical dynamics, but the concept of excursions itself is general and could be used in quantum settings as well. Finally, exploring the role of excursions in generic fluctuation theorems and uncertainty relations, as well as inference of counting observables, is subject to further work.
@@ Acknowledgements.
GF acknowledges useful discussions with Felipe Pereira-Alves, Felipe D. Picoli and Jiheng Duan. This research is primarily supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES) under Award No. DE-SC0025516. PH was supported by the project INTER/FNRS/20/15074473 funded by F.R.S.-FNRS (Belgium) and FNR (Luxembourg).